Determine the Laplace transform of the given generalized function. 78(t-3) ${78(t-3)}(s) =
Answer #1

we are given a problem on Laplace transform.

we first take a look on the Laplace transform rule of delta function below:

L\left \{ \delta (t-a) \right \}=e^{-sa}

Where , a is positive real number.

Now the given problem is,

L\left \{7 \delta (t-3) \right \}

Using the Laplace transform rule given above,

L\left \{7 \delta (t-3) \right \}

=7L\left \{\delta (t-3) \right \}

=7e^{-3s}

therefore the Laplace transform of the given function is 7e-3s .

Thank you for this beautiful question.